### v1

@@ Abstract
In the present work, we investigate the production of $\Xi(1530)$ in the $K^{-}p\to K^{+}\Xi(1530)^{-}$ and $K^{-}p\to K^{0}\Xi(1530)^{0}$ processes by utilizing an effective Lagrangian approach. In order to fit the cross sections for both processes, we include nine $\Lambda/\Sigma$ hyperons and their resonances in the $s$- and $u$-channel processes. Due to the inconsistency of the measured cross sections for $K^{-}p\to K^{+}\Xi(1530)^{-}$ in the range $\sqrt{s}=[2.087,2.168]\ \mathrm{GeV}$, we perform the fitting with all the experimental data having the same weight (model A) and different weight (model B) in the present work. Based on a comparative analysis of the fit results in two models, we conclude that Model A demonstrates a better global agreement with the experimental data than Model B. In addition to the cross sections, the individual contributions from different intermediate states to the cross sections are also estimated . Our results indicate that the cross section resulted from $\Sigma(1193)$ intermediate process is predominant. Moreover, the different cross sections are predicted for both the charged and neutral processes with several typical center-of-mass energies, which could be tested by further experimental measurements at J-PARC in the future.
@@ I Introduction
$\Xi$ hyperons, composed of one up or down quark and two strange quarks, are particularly interesting members of the baryon family because they provide insights into the dynamics of the strong force in environments with high strange quark content. Due to this characteristic, the production of $\Xi$ hyperons is challenging, and experimental observations are limited compared to $\Lambda$ and $\Sigma$ hyperons, which contain only one strange quark. A review of the Particle Data Group (PDG) reveals that only the ground $S$-wave states, $\Xi(1315)^{0,-}$ with $J^{P}=1/2^{+}$ and $\Xi(1530)$ with $J^{P}=3/2^{+}$, are well-established with four-star ratings ParticleDataGroup:2024cfk . Additionally, there are four $\Xi$ resonances with three-star ratings: $\Xi(1690)$, $\Xi(1820)$, $\Xi(1950)$, and $\Xi(2030)$. Of these, the $J^{P}$ quantum numbers of $\Xi(1820)$ have been determined to be $3/2^{-}$, and the spin of $\Xi(2030)$ has been measured to be $5/2$. The remaining five $\Xi$ resonances, $\Xi(1620)$, $\Xi(2120)$, $\Xi(2250)$, $\Xi(2370)$, and $\Xi(2500)$, are only rated as two- or one-star in the PDG ParticleDataGroup:2024cfk , indicating that their existence is only considered to be fairly or even poorly established.
@@
Experimental knowledge of the $\Xi$ hyperon spectrum is severely limited. As highlighted by the Particle Data Group review, a significant portion of the current data on $\Xi$ hyperons derives from low-statistics experiments conducted in the 1960s-1980s with $K^{-}$ beams, and in the 1980s and 1990s with hyperon ($\Sigma^{-}$, $\Xi^{-}$) beams ParticleDataGroup:2024cfk . Despite these limitations, advancements have emerged from collider experiments. Several excited $\Xi$ hyperons have been identified in the decay products of charmed baryons. Notably, the Belle Collaboration detected the $\Xi(1690)$ hyperon in the $\Sigma^{+}K^{-}$ invariant mass from $\Lambda_{c}^{+}\to\Sigma^{+}K^{-}K^{+}$ decays Lyth:2001nq , subsequently confirmed by BaBar in the $\Xi^{-}\pi^{+}$ channel from $\Lambda_{c}^{+}\to\Xi^{-}\pi^{+}K^{+}$ BaBar:2008myc . Belle Collaboration also made the initial observation of $\Xi(1620)^{0}$ decaying to $\Xi^{-}\pi^{+}$ through $\Xi_{c}^{+}\to\Xi^{-}\pi^{+}\pi^{+}$ decays Belle:2018lws . Furthermore, the CLAS Collaboration has contributed through the measurement of $\Xi(1321)^{-}$ hyperon photoproduction via the reactions $\gamma p\rightarrow K^{+}K^{+}\Xi^{-}$ and $\gamma p\rightarrow K^{+}K^{+}\pi^{-}\Xi^{0}$ CLAS:2004gjf ; Price:2004hr , enabling them to obtain a substantial sample of $\Xi(1321)^{-}$. Using a subsequent, high-statistics dataset for the reactions $\gamma p\rightarrow K^{+}K^{+}(X)$ and $\gamma p\rightarrow K^{+}K^{+}\pi^{-}(X)$ Guo:2007dw , the CLAS collaboration measured a mass splitting for the ground state $(\Xi^{-},\Xi^{0})$ doublet, which is $(5.4\pm 1.8)$ MeV  Guo:2007dw .
@@
On the theoretical side, comprehensive investigations of the $\Xi$ spectrum have been undertaken using various approaches. For instance, the spectrum and decay properties of $\Xi$ hyperons have been estimated within the constituent quark model, including relativistic Capstick:1986ter and non-relativistic versions with harmonic confinement potentials Chao:1980em ; Glozman:1995fu ; Xiao:2013xi or linear plus Coulomb potentials Pervin:2007wa . Furthermore, Ref. Bijker:2000gq explores the spectrum and decay properties of $\Xi$ hyperons using a string-like model, where radial excitations are interpreted as rotations and vibrations of the strings. Masses of negative-parity $\Xi$ hyperons have also been analyzed in large $N_{c}$ QCD, considering corrections to order $1/N_{c}$ and first-order SU(3) symmetry breaking Schat:2001xr . Additionally, the mass spectrum and magnetic moments of $\Xi$ baryon resonances have been investigated within the bound-state framework of the Skyrme model Oh:2007cr .
@@
From the perspective of the productions of $\Xi$ states, the two-body process $K^{-}N\to K^{+}\Xi$ is the only feasible binary reaction, while other projectiles such as photons, pions, protons, invariably lead to multi-body final states involving three or more particles. The experimental investigations of $K^{-}p$ scattering began nearly sixty years ago. For example, in Ref. Berge:1966zz the productions of $\Xi^{-}$ and $\Xi^{0}$ hyperons in $K^{-}p$ interactions between 1.05 and 1.7 GeV were measured, and the cross sections were reported to be about $150~{}\mathrm{\mu b}$ for $\Xi^{-}K^{+}$ production and $100~{}\mathrm{\mu b}$ for $\Xi^{0}K^{0}$ and $\Xi\pi K$ productions, in addition to which the production of $\Xi^{\ast}(1530)$ near threshold was also observed. Meanwhile, the $\Xi$-hyperon production in the $K^{-}p$ interaction at $2.24~{}\mathrm{GeV}$ London:1966zz and 3.5 GeV Birmingham-Glasgow-LondonIC-Oxford-Rutherford:1966onr were measured. After that the production cross sections and differential cross sections of $\Xi(1314)$ and $\Xi(1530)$ in the $K^{-}p$ scattering were further reported at different incident kaon beam momenta till to 1977 Trippe:1967wat ; Trower:1968zz ; Burgun:1968ice ; Dauber:1969hg ; SABRE:1971pzp ; deBellefon:1972rjq ; Carlson:1973td ; Griselin:1975pa ; Briefel:1977bp . It is worth noting that the High-Energy Reactions Analysis Group in CERN Flaminio:1979iz presented a lot of cross-sectional experimental data about $\Xi(1314)^{0,-}$ and $\Xi(1530)^{0,-}$ in the $Kp$ scattering processes in 1979, which provides experimental conditions for our study on the production of $\Xi(1530)$.
@@
In the early theoretical researches, there were only a few work progresses on the production of $\Xi$ hyperons due to the limitation of experimental data. In Ref.  Agarwal:1971kb , the authors explained the backward production features of the reaction $K^{-}p\rightarrow\Xi^{-}K^{+}$ by using the two-meson-exchange peripheral model, in which the cross sections and the angular distributions for the reaction $K^{-}p\rightarrow\Xi^{-}K^{+}$ could be well described by using the two-meson-exchange box diagrams. Within the model based on Regge theory, duality and SU(3) symmetry combined with the known data, the authors in Ref. Mir:1981qj investigated the reaction $\bar{K}N\rightarrow\Xi K$ at high energy.
@@
In 2011, the production of $\Xi(1314)$ in the $K^{-}N$ scattering process was investigated in Ref. Sharov:2011xq , where the cross sections and differential cross sections for the $K^{-}N\to K\Xi(1314)$ were fitted within the Phenomenological model. By utilizing the effective Lagrangian approach, the $s$- and $u$-channels exchange by $\Lambda$, $\Sigma$, and their resonances were taken into considerations, and they found that the cross sections and differential cross section could be well described. Similarly, in Ref. Jackson:2015dva , the authors also investigated the process $\bar{K}N\rightarrow\Xi K$ in the effective Lagrangian method. It is worth noting that in addition to the contributions of $s$- and $u$-channels, a contact diagram was also included, and their estimations showed that the dominant contribution came from the contact diagram, while the $s$- and $u$-channels contributions were only about $20\%$ of that of the contact term.
@@
Besides the production of $\Xi(1314)$ in the $K^{-}p$ scattering process, we also notice that there are some data related to cross sections for $K^{-}p\to K^{+}\Xi(1530)^{-}$ and $K^{-}p\to K^{0}\Xi(1530)^{0}$ in Ref, Berge:1966zz ; Dauber:1969hg ; Briefel:1977bp ; Flaminio:1979iz . In addition, the extension project has been proposed in the J-PARC hadron experimental facility, where the high energy kaon beam with high quality can be available. With the kaon beam, the production of $\Xi$-hyperon, including $\Xi(1530)$, will be possible. Thus, we propose to investigate the processes $K^{-}p\to K^{+}\Xi(1530)^{-}$ and $K^{-}p\to K^{0}\Xi(1530)^{0}$ simultaneously in the present work. The model parameters for both the charged and neutral processes could be determined by fitting within the Phenomenological model, and then the cross sections could be estimated with these parameters. In addition, the predicted differential cross sections for $K^{-}p\to K\Xi(1530)$ could be further tested by the experimental measurements at J-PARC in the future.
@@
This work is organized as follows. After the Introduction, the effective approach employed to investigate the processes $K^{-}p\to K\Xi(1530)$ are presented. In Section III, the numerical results and related discussions of the cross sections and the differential cross sections for the relevant processes are presented. The last section is devoted to a summary.
II $\Xi(1530)$ production in the $Kp$ scattering process
@@
Figure 1: Diagrams contributing to the process of $K^{-}p\rightarrow K^{+}\Xi(1530)^{-}$, in which (a) and (b) correspond to the $s$- and $u$-channel contributions, respectively.
@@
In the present work, we employ the effective Lagrangian approach to describe the production of $\Xi(1530)$ in the $K^{-}p$ scattering process. The diagrams contributing to $K^{-}p\to K^{+}\Xi(1530)^{-}$ are presented in Fig. 1, where the $\Lambda/\Sigma$ hyperons and their resonances could serve as intermediate states. In the present calculations, we refer $Y_{J^{P}}$ to the $\Lambda/\Sigma$ hyperon and their resonances with spin $J$ and parity $P$, and the effective Lagrangians for $pY_{1/2^{\pm}}K$ read Jackson:2015dva ,
@@
with the upper and lower symbols in the curly bracket to be the positive and negative parity of $Y$ hyperon, respectively. For the $pY_{3/2^{\pm}}K$ coupling, the effective Lagrangians are Jackson:2015dva ,
@@
In addition, for the hyperon with $J^{P}=5/2^{\pm}$ and $J^{P}=7/2^{\pm}$, the relevant effective Lagrangians read Jackson:2015dva ,
@@
$\mathcal{L}_{pY_{7/2^{\pm}}K}$ $=$ $(-1)\frac{f_{pY_{7/2^{\pm}}K}}{m^{3}_{\pi}}\bar{p}\left\{\begin{array}[]{c}1\\ \gamma_{5}\end{array}\right\}Y^{\mu\nu\rho}_{7/2^{\pm}}\partial_{\mu}\partial_{\nu}\partial_{\rho}K+H.c,\quad$ (12)
@@
$\mathcal{L}_{\Xi^{\prime}Y_{5/2^{\pm}}K}$ $=$ $\frac{f_{\Xi^{\prime}Y_{5/2^{\pm}}K}}{{m_{\pi}}}\bar{\Xi}^{\prime\mu}\left\{\begin{array}[]{c}1\\ \gamma_{5}\end{array}\right\}Y_{5/2^{\pm}}^{\mu\nu}\partial_{\nu}K+H.c,$ (21)
$\mathcal{L}_{\Xi^{\prime}Y_{7/2^{\pm}}K}$ $=$ $(-i)\frac{f_{\Xi^{\prime}Y_{7/2^{\pm}}K}}{{m_{\pi^{2}}}}\bar{\Xi}^{\prime\mu}\left\{\begin{array}[]{c}\gamma_{5}\\ 1\end{array}\right\}Y_{7/2^{\pm}}^{\mu\nu\rho}\partial_{\nu}\partial_{\rho}K+H.c,\quad$ (24)
respectively. Hereafter the $\Xi^{\prime}$ refers to $\Xi(1530)$. In addition, the $\Lambda/\Sigma$ hyperon and their resonances ParticleDataGroup:2024cfk involved in the present estimations are collected in Table 1. With the above effective Lagrangians, one can obtain the amplitudes of the $K^{-}p\to K^{+}\Xi(1530)^{-}$ process, which are,
@@
where the superscripts $s$, $u$ correspond to $s$- and $u$-channel, respectively, and the subscript $Y_{J^{P}}$ corresponds to hyperon involved in the processes. In the above amplitudes, the $\mathcal{S}^{1/2}(k_{i},m_{i},\Gamma_{i})$ refers to the propagator corresponding to the baryon with spin $1/2$, which is Jackson:2015dva ,
$\mathcal{S}^{1/2}(k_{i},m_{i},\Gamma_{i})=\frac{/\!\!\!k_{i}+m_{i}}{k^{2}_{i}-m^{2}_{i}+im_{i}\Gamma_{i}},$ (26)
@@
$\mathcal{S}^{3/2}_{\mu\nu}(k_{i},m_{i},\Gamma_{i})$ $=$ $\frac{/\!\!\!k_{i}+m_{i}}{k^{2}_{i}-m^{2}_{i}+im_{i}\Gamma_{i}}\Big{(}-g^{\mu\nu}+\frac{1}{3}\gamma^{\mu}\gamma^{\nu}$ (27)
@@
$\mathcal{S}^{5/2}_{\mu_{1}\mu_{2}\nu_{1}\nu_{2}}(k_{i},m_{i},\Gamma_{i})$ $=$ $\frac{/\!\!\!k_{i}+\sqrt{s}}{10(k^{2}_{i}-m^{2}_{i}+im_{i}\Gamma_{i})}$ (28)
$\times$ $\Big{(}5P_{\mu_{1}\mu_{2}}P_{\nu_{1}\nu_{2}}-2P_{\mu_{1}\nu_{1}}P_{\mu_{2}\nu_{2}}$
$+$ $5P_{\mu_{1}\nu_{2}}P_{\nu_{1}\mu_{2}}+P_{\mu_{1}\rho}\gamma^{\rho}\gamma^{\sigma}P_{\sigma\mu_{2}}P_{\nu_{1}\nu_{2}}$
$+$ $P_{\nu_{1}\rho}\gamma^{\rho}\gamma^{\sigma}P_{\sigma\nu_{2}}P_{\mu_{1}\mu_{2}}+P_{\mu_{1}\rho}\gamma^{\rho}\gamma^{\sigma}P_{\sigma\nu_{2}}P_{\nu_{1}\mu_{2}}$
$+$ $P_{\nu_{1}\rho}\gamma^{\rho}\gamma^{\sigma}P_{\rho\mu_{2}}P_{\mu_{1}\nu_{2}}\Big{)},$
@@
$P_{\mu\nu}=-g_{\mu\nu}+{q_{\mu}q_{\nu}}/{s}.$ (29)
In addition, two form factors are introduced to depict the two relevant vertices, and both form factors are in the same form, which is
@@
where $\Lambda_{i}$ stands for the cutoff constant. In the present estimations, we adopt two different values of $\Lambda_{i}$, which is $\Lambda_{a}$ for low-spin propagator ($J^{P}\leq 3/2$) and $\Lambda_{b}$ for high-spin propagator ($J\geq 5/2$) as in Ref. Sharov:2011xq . In addition, $\vec{q_{i}}$ and $\vec{q_{f}}$ stand for three-momentum of the initial $K^{-}$ and the final $K^{+}$ in the center-of-mass system, respectively, which read,
@@
Table 1: The resonance parameters of the involved $\Lambda$ and $\Sigma$ hyperon and their resonances.

### v2

@@ Abstract
In the present work, we examine the production of $\Xi(1530)$ in the $K^{-}p\to K^{+}\Xi(1530)^{-}$ and $K^{-}p\to K^{0}\Xi(1530)^{0}$ reactions utilizing an effective Lagrangian approach. To accurately fit the cross sections for both processes, we include nine $\Lambda$ and $\Sigma$ hyperons and their resonances in both $s$- and $u$-channel processes. Considering the discrepancy of the measured cross sections for $K^{-}p\to K^{+}\Xi(1530)^{-}$ within the range $\sqrt{s}=[2.087,2.168]\ \mathrm{GeV}$, we employ two distinct fitting strategies: a uniform weighting scheme (model A) and a different weighting approach (model B). A comparative analysis suggests that model A yields a superior global agreement with experimental data compared to model B. Beyond fitting the cross sections, we also estimate the individual contributions from various intermediate states. Our results reveal that the cross section arising from the $\Sigma(1193)$ intermediate process is dominant. Furthermore, we predict different cross sections for $K^{-}p\to K^{+}\Xi(1530)^{-}$ and $K^{-}p\to K^{0}\Xi(1530)^{0}$ at several representative center-of-mass energies, providing testable predictions for forthcoming J-PARC experiments.
@@ I Introduction
$\Xi$ hyperons, composed of one up or down quark and two strange quarks, are particularly interesting members of the baryon family because they provide insights into the dynamics of the strong force in environments with high strange quark content. Due to this characteristic, the production of $\Xi$ hyperons is challenging, and experimental observations are limited compared to $\Lambda$ and $\Sigma$ hyperons, which contain only one strange quark. A review by the Particle Data Group (PDG) reveals that only the ground $S$-wave states, $\Xi(1315)^{0,-}$ with $J^{P}=1/2^{+}$ and $\Xi(1530)$ with $J^{P}=3/2^{+}$, are well established with four-star ratings ParticleDataGroup:2024cfk . Additionally, there are four $\Xi$ resonances with three-star ratings: $\Xi(1690)$, $\Xi(1820)$, $\Xi(1950)$, and $\Xi(2030)$. Of these, the $J^{P}$ quantum numbers of $\Xi(1820)$ have been determined to be $3/2^{-}$, and the spin of $\Xi(2030)$ has been measured to be $5/2$. The remaining five $\Xi$ resonances, $\Xi(1620)$, $\Xi(2120)$, $\Xi(2250)$, $\Xi(2370)$, and $\Xi(2500)$, are only rated as two or one star ParticleDataGroup:2024cfk , indicating that their existence is only considered to be fairly or even poorly established.
@@
Experimental knowledge of the $\Xi$ hyperon spectrum is severely limited. As highlighted by the Particle Data Group review, a significant portion of the current data on $\Xi$ hyperons derives from low-statistics experiments conducted in the 1960s through the 1980s with $K^{-}$ beams, and in the 1980s and 1990s with hyperon ($\Sigma^{-}$, $\Xi^{-}$) beams ParticleDataGroup:2024cfk . Despite these limitations, advancements have emerged from collider experiments. Several excited $\Xi$ hyperons have been identified in the decay products of charmed baryons. Notably, the Belle Collaboration detected the $\Xi(1690)$ hyperon in the $\Sigma^{+}K^{-}$ invariant mass from $\Lambda_{c}^{+}\to\Sigma^{+}K^{-}K^{+}$ decays Lyth:2001nq , subsequently confirmed by BABAR in the $\Xi^{-}\pi^{+}$ channel from $\Lambda_{c}^{+}\to\Xi^{-}\pi^{+}K^{+}$ BaBar:2008myc . The Belle Collaboration also made the initial observation of $\Xi(1620)^{0}$ decaying to $\Xi^{-}\pi^{+}$ through $\Xi_{c}^{+}\to\Xi^{-}\pi^{+}\pi^{+}$ decays Belle:2018lws . Furthermore, the CLAS Collaboration has contributed through the measurement of $\Xi(1321)^{-}$ hyperon photoproduction via the reactions $\gamma p\rightarrow K^{+}K^{+}\Xi^{-}$ and $\gamma p\rightarrow K^{+}K^{+}\pi^{-}\Xi^{0}$ CLAS:2004gjf ; Price:2004hr , enabling them to obtain a substantial sample of $\Xi(1321)^{-}$. Using a subsequent, high-statistics dataset for the reactions $\gamma p\rightarrow K^{+}K^{+}(X)$ and $\gamma p\rightarrow K^{+}K^{+}\pi^{-}(X)$ Guo:2007dw , the CLAS collaboration measured a mass splitting for the ground state $(\Xi^{-},\Xi^{0})$ doublet, which is $(5.4\pm 1.8)$ MeV  Guo:2007dw .
@@
On the theoretical side, comprehensive investigations of the $\Xi$ spectrum have been undertaken using various approaches. For instance, the spectrum and decay properties of $\Xi$ hyperons have been estimated within the constituent quark model, including relativistic Capstick:1986ter and nonrelativistic versions with harmonic confinement potentials Chao:1980em ; Glozman:1995fu ; Xiao:2013xi or linear plus Coulomb potentials Pervin:2007wa . Furthermore, Ref. Bijker:2000gq explores the spectrum and decay properties of $\Xi$ hyperons using a string like model, where radial excitations are interpreted as rotations and vibrations of the strings. Masses of negative-parity $\Xi$ hyperons have also been analyzed in large-$N_{c}$ QCD, considering corrections to order $1/N_{c}$ and first-order SU(3) symmetry breaking Schat:2001xr . Additionally, the mass spectrum and magnetic moments of $\Xi$ baryon resonances have been investigated within the bound-state framework of the Skyrme model Oh:2007cr .
@@
From the perspective of the productions of $\Xi$ states, the two-body process $K^{-}N\to K^{+}\Xi$ is the only feasible binary reaction, while other projectiles such as photons, pions, and protons, invariably lead to multibody final states involving three or more particles. The experimental investigations of $K^{-}p$ scattering began nearly 60 years ago. For example, in Ref. Berge:1966zz the productions of $\Xi^{-}$ and $\Xi^{0}$ hyperons in $K^{-}p$ interactions between 1.05 and 1.7 GeV were measured, and the cross sections were reported to be about $150~\mathrm{\mu b}$ for $\Xi^{-}K^{+}$ production and $100~\mathrm{\mu b}$ for $\Xi^{0}K^{0}$ and $\Xi\pi K$ productions, in addition to which the production of $\Xi^{\ast}(1530)$ near threshold was also observed. Meanwhile, the $\Xi$-hyperon production in the $K^{-}p$ interaction at $2.24~\mathrm{GeV}$ London:1966zz and 3.5 GeV Birmingham-Glasgow-LondonIC-Oxford-Rutherford:1966onr was measured. After that the production cross sections and differential cross sections of $\Xi(1314)$ and $\Xi(1530)$ in the $K^{-}p$ scattering were further reported at different incident kaon beam momenta until to 1977 Trippe:1967wat ; Trower:1968zz ; Burgun:1968ice ; Dauber:1969hg ; SABRE:1971pzp ; deBellefon:1972rjq ; Carlson:1973td ; Griselin:1975pa ; Briefel:1977bp . It is worth noting that the High-Energy Reactions Analysis Group in CERN Flaminio:1979iz presented a lot of cross-sectional experimental data about $\Xi(1314)^{0,-}$ and $\Xi(1530)^{0,-}$ in the $Kp$ scattering processes in 1979, which provides experimental conditions for our study on the production of $\Xi(1530)$.
@@
In early theoretical research, only a few works made progress on the production of $\Xi$ hyperons due to the limitation of experimental data. In Ref.  Agarwal:1971kb , the authors explained the backward production features of the reaction $K^{-}p\rightarrow\Xi^{-}K^{+}$ by using the two-meson-exchange peripheral model, in which the cross sections and the angular distributions for the reaction $K^{-}p\rightarrow\Xi^{-}K^{+}$ could be well described by using the two-meson-exchange box diagrams. Within the model based on Regge theory, duality and SU(3) symmetry combined with the known data, the authors in Ref. Mir:1981qj investigated the reaction $\bar{K}N\rightarrow\Xi K$ at high energy.
@@
In 2011, the production of $\Xi(1314)$ in the $K^{-}N$ scattering process was investigated Sharov:2011xq , where the cross sections and differential cross sections for the $K^{-}N\to K\Xi(1314)$ reaction were fitted within the phenomenological model. By utilizing the effective Lagrangian approach, the $s$- and $u$-channel exchange by $\Lambda$, $\Sigma$, and their resonances were taken into consideration, and they found that the cross sections and differential cross section could be well described. Similarly, in Ref. Jackson:2015dva , the authors also investigated the process $\bar{K}N\rightarrow\Xi K$ in the effective Lagrangian method. It is worth noting that in addition to the contributions of $s$ and $u$ channel, a contact diagram was also included, and their estimations showed that the dominant contribution came from the contact diagram, while the $s$- and $u$-channel contributions were only about $20\%$ of that of the contact term. Moreover, based on a hybrid Regge-plus-resonance model, the authors of Ref. Kim:2023jij also studied the process $\bar{K}N\rightarrow\Xi K$. The estimations showed that the bump structures in the total cross sections could be described by the $s$-channel hyperon resonances, while the dominant contributions come from $\Lambda(2100)$ and $\Sigma(2030)$ resonances.
@@
In addition to the production of $\Xi(1314)$ in the $K^{-}p$ scattering process, we also notice that there have been some data related to cross sections for $K^{-}p\to K^{+}\Xi(1530)^{-}$ and $K^{-}p\to K^{0}\Xi(1530)^{0}$ in Refs Berge:1966zz ; Dauber:1969hg ; Briefel:1977bp ; Flaminio:1979iz . Moreover, the extension project has been proposed in the J-PARC hadron experimental facility, where the high energy kaon beam with high quality can be available. With the kaon beam, the production of $\Xi$ hyperon, including $\Xi(1530)$, will be possible. Thus, we propose to investigate the processes $K^{-}p\to K^{+}\Xi(1530)^{-}$ and $K^{-}p\to K^{0}\Xi(1530)^{0}$ in the present work. By fitting the cross sections of two processes, the model parameters could be determined, and then the differential cross sections are predicted, which can further be tested by the experimental measurements at J-PARC in the future.
@@
This work is organized as follows. After the Introduction, the effective Lagrangian approach employed to investigate the processes $K^{-}p\to K\Xi(1530)$ are presented. In Sec.III, the numerical results and relevant discussions of the cross sections and the differential cross sections for the considered processes are presented. The last section is devoted to a short summary.
II $\Xi(1530)$ production in the $K^{-}p$ scattering process
@@
Figure 1: Diagrams contributing to the process of $K^{-}p\rightarrow K^{+}\Xi(1530)^{-}$, corresponding to the (a) $s$- and (b) $u$-channel contributions, respectively.
@@
In the present work, we employ an effective Lagrangian approach to describe the interactions of hadrons in the $K^{-}p\to K\Xi(1530)$ process. The diagrams contributing to $K^{-}p\to K^{+}\Xi(1530)^{-}$ are presented in Fig. 1, where $\Lambda$ and $\Sigma$ hyperons and their resonances could serve as intermediate states. In the following, we assign the notation $Y_{J^{P}}$ to represent $\Lambda$ and $\Sigma$ hyperons and their resonance states characterized by spin quantum number $J$ and parity $P$. In the present calculations, we only take the lower spin states with $J\leq 5/2$ into consideration, as these states are expected to dominate the physical processes occurring near the production threshold region. The effective Lagrangians for $pY_{1/2^{\pm}}K$ read Jackson:2015dva ,
@@
where the upper and lower symbols in the curly bracket are the positive and negative parity of the $Y$ hyperon, respectively. For $pY_{3/2^{\pm}}K$ and $pY_{5/2^{\pm}}K$ couplings, the effective Lagrangians are Jackson:2015dva ,
@@
$\mathcal{L}_{\Xi^{\prime}Y_{5/2^{\pm}}K}$ $=$ $\frac{f_{\Xi^{\prime}Y_{5/2^{\pm}}K}}{{m_{\pi}}}\bar{\Xi}^{\prime\mu}\left\{\begin{array}[]{c}1\\ \gamma_{5}\end{array}\right\}Y_{5/2^{\pm}}^{\mu\nu}\partial_{\nu}K+H.c,\quad$ (18)
respectively. Hereafter the $\Xi^{\prime}$ refers to $\Xi(1530)$. The details of the involved $\Lambda$ and $\Sigma$ resonances will be discussed in the following sections. With the above effective Lagrangians, one can obtain the amplitudes of the $K^{-}p\to K^{+}\Xi(1530)^{-}$ process, which are,
@@
where the superscripts $s$ and $u$ correspond to $s$ and $u$ channels, respectively, and the subscript $Y_{J^{P}}$ corresponds to the hyperon involved in the processes. In the above amplitudes, the $\mathcal{S}^{1/2}(k_{i},m_{i},\Gamma_{i})$ refers to the propagator corresponding to the baryon with spin $1/2$, which is Jackson:2015dva ,
$\mathcal{S}^{1/2}(k_{i},m_{i},\Gamma_{i})=\frac{/\penalty 50\!\!\!k_{i}+m_{i}}{k^{2}_{i}-m^{2}_{i}+im_{i}\Gamma_{i}},$ (20)
@@
$\mathcal{S}^{3/2}_{\mu\nu}(k_{i},m_{i},\Gamma_{i})$ $=$ $\frac{/\penalty 50\!\!\!k_{i}+m_{i}}{k^{2}_{i}-m^{2}_{i}+im_{i}\Gamma_{i}}\Big(-g^{\mu\nu}+\frac{1}{3}\gamma^{\mu}\gamma^{\nu}$ (21)
@@
$\mathcal{S}^{5/2}_{\mu_{1}\mu_{2}\nu_{1}\nu_{2}}(k_{i},m_{i},\Gamma_{i})$ $=$ $\frac{/\penalty 50\!\!\!k_{i}+\sqrt{s}}{10(k^{2}_{i}-m^{2}_{i}+im_{i}\Gamma_{i})}$ (22)
$\times$ $\Big(5\tilde{g}_{\mu_{1}\mu_{2}}\tilde{g}_{\nu_{1}\nu_{2}}-2\tilde{g}_{\mu_{1}\nu_{1}}\tilde{g}_{\mu_{2}\nu_{2}}$
$+$ $5\tilde{g}_{\mu_{1}\nu_{2}}\tilde{g}_{\nu_{1}\mu_{2}}+\tilde{g}_{\mu_{1}\rho}\gamma^{\rho}\gamma^{\sigma}\tilde{g}_{\sigma\mu_{2}}\tilde{g}_{\nu_{1}\nu_{2}}$
$+$ $\tilde{g}_{\nu_{1}\rho}\gamma^{\rho}\gamma^{\sigma}\tilde{g}_{\sigma\nu_{2}}\tilde{g}_{\mu_{1}\mu_{2}}+\tilde{g}_{\mu_{1}\rho}\gamma^{\rho}\gamma^{\sigma}\tilde{g}_{\sigma\nu_{2}}\tilde{g}_{\nu_{1}\mu_{2}}$
$+$ $\tilde{g}_{\nu_{1}\rho}\gamma^{\rho}\gamma^{\sigma}\tilde{g}_{\rho\mu_{2}}\tilde{g}_{\mu_{1}\nu_{2}}\Big),$
@@
$\tilde{g}_{\mu\nu}=-g_{\mu\nu}+{q_{\mu}q_{\nu}}/{s}.$ (23)
In addition, two form factors are introduced to depict the two relevant vertices, and both form factors are in the form,
@@
where $\Lambda_{i}$ stands for the cutoff constant. In the present estimations, we adopt two different values of $\Lambda_{i}$, which is $\Lambda_{a}$ for low-spin propagator ($J^{P}\leq 3/2$) and $\Lambda_{b}$ for high-spin propagator ($J\geq 5/2$) as in Ref. Sharov:2011xq . In addition, $\vec{q_{i}}$ and $\vec{q_{f}}$ stand for three momenta of the initial $K^{-}$ and the final $K^{+}$ in the center-of-mass system, respectively, which read,
@@
Figure 2: Diagrams contributing to the process of $K^{-}p\rightarrow K^{0}\Xi(1530)^{0}$, corresponding to the (a) $s$- and (b) $u$-channel contributions, respectively.
In the same way, one can investigate the $\Xi(1530)^{0}$ production in the $K^{-}p$ scattering process, and the relevant diagrams are shown in Fig. 2. It should be noted that in the $u$ channel, the exchanged hyperons could only be $\Sigma$ and its resonances, while the $\Lambda$ hyperon and its resonances are forbidden due to charge conservation. Thus, from the isospin symmetry, the contributions from the $\Sigma$ hyperon and its resonances are the same in both $K^{-}p\to K^{+}\Xi(1530)^{-}$ and $K^{-}p\to K^{0}\Xi(1530)^{0}$ processes, while those from the $\Lambda$ hyperon and its resonances are different in both processes.
III NUMERICAL RESULTS AND DISCUSSIONS
Table 1: The resonance parameters of the involved $\Lambda$ and $\Sigma$ hyperons and their resonances.